OpenStudy (anonymous):

Need Help Fasttt! Medalll What is the eqaution of the line that passes through the points (1,2/3) and (-1,-1)?

OpenStudy (anonymous):

@iGreen can you help plzzzzzz

OpenStudy (anonymous):

@JungHyunRan

OpenStudy (anonymous):

Equation of the point passing through 2 points is given by: \(\Large \frac{y-y_1}{y_2-y_1} = \frac{x-x_1}{x_2-x_1}\) The two points are: \(\Large (1,\frac{2}{3})\) and \(\Large(-1, -1)\) plugging in these points: \(\huge \frac{y-\frac{2}{3}}{-1-\frac{2}{3}}=\frac{x-1}{-1-1}\\\huge (y-\frac{2}{3})(-2)=(x-1)(-\frac{5}{3})\\\huge -2y+\frac{4}{3}=\frac{-5}{3}x+\frac{5}{3}\\\huge2y -\frac{5}{3}x+\frac{5}{3}-\frac{4}{3}=0\\\huge6y-5x+1=0\\\huge y=\frac{5x-1}{6}\)

OpenStudy (anonymous):

What is the equation of the line passing through the points (–2, 2) and (3, –1)?

OpenStudy (anonymous):

@JungHyunRan

OpenStudy (anonymous):

it is similar

OpenStudy (anonymous):

\(\Large x_1=-2 , y_1=2\ and\ x_2=3 , y_2=-1\) So, \(\huge \frac{y-y_1}{y_2-y_1} = \frac{x-x_1}{x_2-x_1}\\\huge \frac{y-2}{-1-2}=\frac{x+2}{3+2}\\\Large 5(y-2)=-3(x+2)\\\Large 5y-10=-3x-6\\\Large y=\frac{-3x+4}{5}\)